A car carrying a driver and a front-seat passenger of mass 60 kg initially is moving with a speed 18 m/s. The car is then slowed down to rest by an air bag in a time period of 0.4 s. Find the average force in N on the front-seat passenger during this time period. First calculate the impulse and then the average force. Do not use a comma to separate every three-digits.


1
Expert's answer
2021-11-10T10:10:43-0500

According to the second Newton's law, the impulse is given as follows:


"J = m(v_2 - v_1)"

where "m = 60kg, v_1 = 18m\/s, v_2 = 0m\/s". Thus, obtain:


"J = 60\cdot (0 - 18)=-1080kg\cdot m/s"

The average force is then:


"F = \dfrac{J}{\Delta t}"

where "\Delta t = 0.4s". Thus, the force is:


"F = \dfrac{-1080kg\cdot m/s}{0.4s} = 2700N"

Answer. 2700 N.